The equation of the plane through the points (0,1,8),(1,1,11) and (1,6,0) is:



z= Blank 1. Calculate the answer by read surrounding text.


+ Blank 2. Calculate the answer by read surrounding text.


x+ Blank 3. Calculate the answer by read surrounding text.


y.




1
Expert's answer
2022-03-17T18:30:45-0400

"A(0,1,8), B(1,1,11), C(1,6,0)\\
\begin{vmatrix}
 x-0 & y-1&z-8 \\
 1-0 & 1-1&11-8\\
1-0&6-1&0-8
\end{vmatrix}=0\\
\begin{vmatrix}
 x & y-1&z-8 \\
 1 & 0&3\\
1&5&-8
\end{vmatrix}=0\\
x\cdot \begin{vmatrix}
 0 & 3 \\
 5 & -8
\end{vmatrix}-(y-1)\cdot \begin{vmatrix}
 1 & 3 \\
 1 & -8
\end{vmatrix}+\\+(z-8)\cdot \begin{vmatrix}
 1 & 0 \\
 1 & 5
\end{vmatrix}=0\\
(0-15)x-(y-1)(-8-3)+\\+(z-8)(5-0)=0\\
-15x+11(y-1)+5(z-8)=0\\
-15x+11y+5z-51=0"